# Optimal Denial-of-Service Attacks Against Partially-Observable Real-Time Monitoring Systems

### v1

@@ Abstract
In this paper, we investigate the impact of denial-of-service attacks on the status updating of a cyber-physical system with one or more sensors connected to a remote monitor via unreliable channels. We approach the problem from the perspective of an adversary that can strategically jam a subset of the channels. The sources are modeled as Markov chains, and the performance of status updating is measured based on the age of incorrect information at the monitor. Our objective is to derive jamming policies that strike a balance between the degradation of the system’s performance and the conservation of the adversary’s energy. For a single-source scenario, we formulate the problem as a partially-observable Markov decision process, and rigorously prove that the optimal jamming policy is of a threshold form. We then extend the problem to a multi-source scenario. We formulate this problem as a restless multi-armed bandit, and provide a jamming policy based on the Whittle’s index. Our numerical results highlight the performance of our policies compared to baseline policies.
@@ I Introduction
Cyber-physical systems are complex systems that integrate computational algorithms with dynamical processes, thereby unifying the digital and physical worlds seamlessly [1]. These systems use computation, communication, and control to improve efficiency, adaptability, and automation in various modern domains such as smart cities, factories, healthcare, and transportation. Due to their highly dynamic nature, cyber-physical systems require real-time monitoring so that the latest changes in the physical environment is captured. This real-time flow of data enables these systems to quickly respond to the new conditions, ensuring decisions are made based on the most current and relevant information [2, 3, 4, 5, 6, 7]. For example, autonomous vehicles navigating traffic and smart grids managing energy distribution both rely on real-time status updates. Autonomous vehicles must constantly analyze sensory information to make navigation decisions, and adapt to road conditions and traffic patterns. Similarly, smart grids should constantly use sensory information to balance energy supply and demand, and allocate resources optimally. In these cases, real-time decision-making depends on the timeliness and accuracy of status updates.
@@
However, cyber-physical systems are inherently vulnerable to various forms of cyber attacks [8]. Three prevalent types of cyber attacks include deception attacks, eavesdropping attacks, and denial-of-service attacks, as noted in [9]. Deception attacks compromise data integrity, eavesdropping attacks breach confidentiality, and denial-of-service attacks—the focus of our study—disrupt data availability by obstructing network transmissions. Here, we investigate the impact of denial-of-service attacks on the status updating of a cyber-physical system with one or more sensors connected to a remote monitor via unreliable channels. We approach this problem from the perspective of an adversary that can strategically jam the channels. Our objective is to derive jamming policies that strike a balance between the degradation of the system’s performance and the conservation of the adversary’s energy. Determining such policies enable us to gain insight into the potential damage that such attacks can inflict on monitoring systems.
@@
In our study, the sources are modeled as Markov chains, and the performance of status updating is measured based on the age of incorrect information (AoII) at the monitor in a partially-observable environment. The AoII was initially proposed in [10], which takes into account both the change of the source status and the freshness of information. The reader is referred to [11, 5, 12, 13, 14], and the references therein, for more details about this semantic metric and its use for cyber-physical systems. Notably, the works in [10, 5, 12] developed optimal scheduling policies that minimize the AoII in real-time monitoring applications, where it is assumed that the scheduler has a perfect knowledge of the status of the source at each time. However, in practice, specially if multiple sources need to be tracked, the scheduler might not be aware of the status of the sources [11]. In this partially-observable environment, the AoII is not known, and must be estimated.
There exists a body of research on the design of jamming polices in the context of cyber-physical systems [15, 16, 17, 18, 19, 20, 21]. Notably, Zhang et al. [15] obtained the optimal signal-independent jamming policy that maximizes a regulation loss function subject to a blockage frequency constraint in a networked control system, and extended the results to a multi-source scenario. Zhang et al. [16] also derived the optimal signal-independent jamming policy that maximizes a distortion loss function subject to a blockage frequency constraint in a networked estimation system, and extended the results to the case in which there exists a packet-drop-ratio stealthiness constraint. Qin et al. [17] found the optimal signal-independent jamming policy that maximizes a distortion loss function subject to a blockage frequency constraint in a networked estimation system when the channel is subject to packet loss even in the absence of attacks, and extended the results to a multi-source scenario. Zhang et al. [18] obtained the optimal signal-independent jamming policy that maximizes a distortion loss function subject to a jamming power constraint in a networked estimation system when the channel is ideal in the absence of attacks, and presented their findings for both static and dynamic attacks. Qin et al. [19] extended the results in  [18] to the case in which the channel is subject to packet loss even in the absence of attacks. Gan et al. [20] found the optimal signal-independent jamming policy that maximizes a distortion loss function subject to a jamming power constraint in a networked estimation system when the sensor is connected to the monitor via a two-hop relay channel. In addition, Zhang et al. [21] obtained the optimal signal-dependent jamming policy that maximizes a regulation loss function subject to a soft jamming power constraint in a networked control system when the channel is fading, and studied different packet detection schemes.
@@
To the best of our knowledge, there are only a few works that have addressed the design of jamming policies with freshness metrics [22, 23, 24, 25]. In particular, the work in [22] formulated the interaction between the adversary and the transmitter as a static game with an age of information (AoI) utility function, and obtained the optimal jamming power level when the channel is modeled as a $M$/$G$/$1$/$1$ queue. The work in [23] formulated the interaction between the adversary and the transmitter as a dynamic game with an AoI utility function, and derived the optimal jamming time distribution when the channel is reliable. The work in [24], focused on coordinating multiple sensors’ channel access behavior in a distributed way to minimize the average AoI under a jamming attack, and studied the channel access policies when the channel is modeled as $M$/$M$/$1$, $M$/$D$/$1$, $D$/$M$/$1$ queues. The work in [25], which presents our preliminary results, considered a single-source scenario and derived the optimal jamming policy of an omniscient adversary that knows the mismatch between the true state and the estimated state at the monitor at each time.
@@ II-A Networked System Model
We consider a binary Markovian source that generates and instantaneously transmits status updates in the format of data packets to a remote monitor over an independent and identically distributed (i.i.d.) channel. Time is discretized and normalized to the time slot duration. For the source, the probability of transitioning from one state to another is $r$. We assume that $r<1/2$. Let $c(t)$ denote the quality of the channel during regular operation at time $t$ such that $c(t)=1$ if the transmitted packet is successfully decoded by the receiver in the absence of any attack, and $c(t)=0$ if a packet loss occurs. The probability of $c(t)=1$ is equal to $p$. We assume that the receiver instantaneously sends an acknowledgement (ACK) to the transmitter when a packet is correctly decoded, and sends a negative-acknowledgment (NACK) otherwise.
An adversary, acting as an agent within our system model, attempts to jam the channel. Let $a(t)$ denote the action of the adversary at time $t$ such that $a(t)=1$ if the adversary attempts to jam the channel, and $a(t)=0$ otherwise. When the adversary decides to jam the channel, the probability that it succeeds is equal to $q$. We assume that the adversary can intercept the ACK and NACK messages at each time.
@@ II-B Performance Index and It’s Dynamics
Note that, at time $t+1$, if the packet is successively transmitted, then $g(t+1)=t+1$. Therefore, at time $t+1$, the EAoII equals to $s(t+1)=s_{t+1-g(t+1)}=s_{0}$. This occurs if $c(t+1)=1$ and $a(t)=0$; or $c(t+1)=1$ and $a(t)=1$ but the adversary does not succeed in jamming the channel. However, if the packet is not successively transmitted, then $g(t+1)=g(t)$. Therefore, the EAoII will transition to $s(t+1)=s_{t+1-g(t+1)}=s_{t-g(t)+1}$. Following this observation, the transition probabilities for the EAoII are obtained as
$\Pr\big{(}s(t+1)=s_{0}|s(t)=s_{k},a(t)=0\big{)}=p,$
$\Pr\big{(}s(t+1)=s_{k+1}|s(t)=s_{k},a(t)=0\big{)}=1-p,$
$\Pr\big{(}s(t+1)=s_{0}|s(t)=s_{k},a(t)=1\big{)}=p(1-q),$
$\Pr\big{(}s(t+1)=s_{k+1}|s(t)=s_{k},a(t)=1\big{)}=\!-p(1\!-\!q).$ (4)
@@ II-C Problem Statement
We suppose that, when the adversary decides to jam the channel at time $t$, an additional positive cost is incurred due to the energy consumed for jamming. A jamming policy $\phi$ is defined as a sequence of actions $\phi=(a^{\phi}(0),a^{\phi}(1),\ldots)$, where $a^{\phi}(t)=1$ if the adversary decides to jam the channel at time $t$, and $a^{\phi}(t)=0$ otherwise. Let $R(t)=s(t)-\lambda a(t)$ be the reward function of the adversary at time $t$, where $\lambda>0$. Our aim is to find the optimal jamming policy that maximizes the expected average reward. This can be formulated as
@@ II-D Optimal Jamming Policy
In the rest of this section, we provide our theoretical results on the design of the optimal jamming policy. We begin by showing that the optimal solution is a threshold policy. Then, we provide a closed-form expression of the problem of interest, and derive the optimal threshold value as a function of $\lambda$.
The optimization problem in (5) can be viewed as an infinite-horizon average-reward POMDP with the following characteristics:
  • State: The state at time $t$, denoted by $s(t)$, is equal to the EAoII at time $t$.
  • Action: The action at time $t$, denoted by $a(t)$, determines if the channel is attacked (i.e., $a(t)=1$) or not (i.e., $a(t)=0$).
  • Transition Probabilities: The transition probabilities specify the probabilities associated with transitioning from $s(t)$ to $s(t+1)$ given an action $a(t)$.
  • Reward: The instantaneous reward, denoted by $R(t)$, is equal to $s(t)-\lambda a(t)$.
The optimal policy $\phi^{*}$ of the problem in (5) can be obtained by solving the following Bellman optimality equation for each state $s=s(t)$:
$\theta+V(s)=\max_{a\in\{0,1\}}\Big{\{}s-\lambda a+\sum_{s^{\prime}\in\mathbb{S}}\Pr(s\rightarrow s^{\prime}|a)V(s^{\prime})\Big{\}}$ (6)
where $\Pr(s\rightarrow s^{\prime}|a)$ is the transition probability from state $s$ to $s^{\prime}$ given the action $a$, $\theta$ is the optimal value of the problem, $V(s)$ is the differential reward-to-go function, and $\mathbb{S}$ is a set defined as
@@
The optimal jamming policy associated with the problem in (6) and for any given $\lambda$ is an increasing threshold policy. More specifically, there exists $n\in\mathbb{N}$ such that the prescribed action is passive, i.e., $a_{k}=0$, when $s_{k}<s_{n}$, and the prescribed action is active, i.e., $a_{k}=1$, when $s_{k}\geq s_{n}$.
@@ II-E Closed-Form Expression of the Optimal Solution
We have proved that the optimal solution of the problem in (5) is a threshold policy. Nevertheless, we still have to determine the optimal threshold value for any given $\lambda$. To that end, we will first derive a simple closed-form expression for the average reward. Note that we can derive the steady-state form of the problem in (5) for a given threshold jamming policy with parameter $s_{n}$ as
$\underset{n\in\mathbb{N}}{\text{maximize}}\ \overline{s}_{n}-\lambda\overline{a}_{n}$ (16)
where $\overline{s}_{n}$ is the average EAoII, and $\overline{a}_{n}$ is the average active attacking time (AAT) under the threshold jamming policy with parameter $s_{n}$. More specifically, we have
@@
$\overline{a}_{n}$ $=\lim_{T\to\infty}\text{inf}\>\frac{1}{T}\mathbb{E}^{n}\Big{[}\sum_{t=0}^{T-1}a(t)\big{|}s(0),\phi^{t}(n)\Big{]}$ (18)
@@
To compute $\overline{s}_{n}$ and $\overline{a}_{n}$, we will determine the stationary distribution of the discrete-time Markov chain (DTMC) that models the evolution of the EAoII under the threshold policy with parameter $s_{n}$. We first provide our result for a more general DTMC, illustrated in Fig. 1.
@@
$u_{n}(s_{i})=\left\{\begin{array}[]{ll}\frac{(1-q)p}{1-q+(1-p)^{n}}&\text{if}\ i=0,\\[9.0pt] (1\!-\!p)^{i}\frac{(1-q)p}{1-q+(1-p)^{n}}&\text{if}\ 1\leq i\leq n,\\[9.0pt] (1\!-\!p)^{n}(pq+1\!-\!p)^{i-n}&\text{if}\ i\geq n+1.\\[3.0pt] \qquad\qquad\times\frac{(1-q)p}{1-q+(1-p)^{n}}\end{array}\right.$ (23)
@@
$\overline{s}_{n}$ $=\frac{(1-q)p}{1-q+(1-p)^{n}}\bigg{[}\frac{r(1-p)}{p(1-y)(1-z)}$
$\quad+\frac{q(1-p)^{n}}{2pr(1-q)}-\frac{pq(1-p)^{n}(1-r)^{n+2}}{r(1-z)\big{(}1-b(1-r)\big{)}}$
$\quad+\frac{pq(1-p)^{n}(1-2r)^{n+2}}{2r(1-y)\big{(}1-b(1-2r)\big{)}}\bigg{]}$ (25)
where $y=(1-2r)(1-p)$ and $z=(1-r)(1-p)$.
@@
Under a threshold jamming policy $s_{n}$, the average AAT $\overline{a}_{n}$ is equal to
$\overline{a}_{n}=\frac{(1-p)^{n}}{(1-q)+q(1-p)^{n}}.$ (26)
@@
Again, in the steady state, $\overline{a}_{n}=\sum_{k=n}^{\infty}u_{n}(s_{k})$. Hence, exploiting the expression of $u_{n}(\cdot)$ given in Proposition 1, we can obtain the result. ∎
Now that we have derived the steady-state form of the problem in (5), we can search for the threshold value $s_{n}$ that maximizes $\overline{s}_{n}-\lambda\overline{a}_{n}$. The sequence $\lambda(s_{n})$ is defined by the following definition, with its exact expression resembling that in (3).
@@
$\lambda(s_{n})$ is the intersection point between $\overline{s}_{n}-\lambda\overline{a}_{n}$ and $\overline{s}_{n+1}-\lambda\overline{a}_{n+1}$, i.e.,
$\lambda(s_{n})=\frac{\overline{s}_{n+1}-\overline{s}_{n}}{\overline{a}_{n+1}-\overline{a}_{n}}.$ (27)
@@
$\overline{s}_{k}\!-\!\lambda\overline{a}_{k}\geq\overline{s}_{k+1}\!-\!\lambda\overline{a}_{k+1}\ \text{for}\ \lambda\leq\lambda(s_{k}),$ (29)
$\overline{s}_{k}\!-\!\lambda\overline{a}_{k}<\overline{s}_{k+1}\!-\!\lambda\overline{a}_{k+1}\ \text{for}\ \lambda>\lambda(s_{k}),$ (30)
$\overline{s}_{n+1}\!-\!\lambda\overline{a}_{n+1}>\overline{s}_{k}\!-\!\lambda\overline{a}_{k}\ \text{for}\ \lambda>\lambda(s_{n}),\ k\leq n,$ (31)
$\overline{s}_{n+1}\!-\!\lambda\overline{a}_{n+1}\geq\overline{s}_{k}\!-\!\lambda\overline{a}_{k}\ \text{for}\ \lambda\leq\lambda(s_{n\!+\!1}),\ k>n.$ (32)

### v2

@@ Abstract
In this paper, we study the impact of denial-of-service attacks on the status updating of a cyber-physical system with one or more sensors connected to a remote monitor via unreliable channels in partially-observable environment. We approach the problem from the perspective of an adversary that can strategically jam a subset of the channels. The sources are modeled as Markov chains, and the performance of status updating is measured based on the age of incorrect information at the monitor. Our goal is to derive jamming policies that severely degrade the system’s performance while operating within the adversary’s energy constraint. For a single-source scenario, we formulate the problem as a partially-observable Markov decision process, and rigorously prove that the optimal jamming policy is of a threshold form. We then extend the problem to a multi-source scenario. We formulate this problem as a restless multi-armed bandit, and provide a jamming policy based on the Whittle’s index. Our numerical results highlight the performance of our policies compared to baseline policies.
@@ I Introduction
Cyber-physical systems (CPSs) represent a sophisticated fusion of computational algorithms and physical processes, bridging the gap between the digital and real-world environments [1]. These systems leverage advanced computation, communication, and control mechanisms to enhance performance, adaptability, and automation across various sectors. The dynamic nature of CPSs demands continuous real-time monitoring to capture the most recent changes in their operating environments. Such a stream of up-to-date data allows CPSs to swiftly adapt to evolving conditions, ensuring decisions are based on accurate and timely information [2, 3, 4, 5, 6, 7]. Nevertheless, CPSs are inherently susceptible to a wide range of cyber threats and attacks [8]. Three prevalent types of cyber attacks include deception attacks, eavesdropping attacks, and denial-of-service attacks, as noted in [9]. Deception attacks compromise data integrity, eavesdropping attacks breach confidentiality, and denial-of-service attacks—the focus of our study—disrupt data availability by obstructing network transmissions. We here investigate the impact of denial-of-service attacks on the status updating of a CPS with one or more sensors connected to a remote monitor via unreliable channels. We approach this problem from the perspective of an adversary that can strategically jam the channels. Our objective is to derive jamming policies that severely degrade the system’s performance while operating within the adversary’s energy constraint. Determining such policies enable us to gain insight into the potential damage that such attacks can inflict on monitoring systems.
@@
In our study, the sources are modeled as Markov chains, and the performance of status updating is measured based on the age of incorrect information (AoII) at the monitor in a partially-observable environment. The AoII was initially proposed in [10], which takes into account both the change of the source status and the freshness of information. The reader is referred to [11, 5, 12, 13, 14], and the references therein, for more details about this semantic metric and its use for CPSs. Notably, the works in [10, 5, 12] developed optimal scheduling policies that minimize the AoII in real-time monitoring applications, where it is assumed that the scheduler has a perfect knowledge of the status of the source at each time. However, in practice, specially if multiple sources need to be tracked, the scheduler might not be aware of the status of the sources [11]. In this partially-observable environment, the AoII is not known, and must be estimated.
Previous research have explored the designing jamming policies within the context of networked estimation and networked control [15, 16, 17, 18, 19, 20, 21]. For instance, optimal jamming policy maximizing a regulation loss subject to a frequency constraint has been developed in [15]. Packet-drop-ratio stealthiness is considered in [16] where a distortion loss subject to a frequency constraint is optimized using signal-independent jamming policy. [17] considers a channel with inherent packet loss for multi-source scenario and identified an optimal signal-independent jamming policy that maximizes distortion loss subject to a frequency constraint. In [18], static and dynamic attack scenarios are considered and optimal jamming policy that maximizes distortion loss subject to a power constraint is developed, while considering that the channel is ideal in the absence of attacks. Building on [18], a more challenging scenario is considered in [19] where packet loss can occur even without any attacks. In [20], a scenario where sensors communicate with monitors via two-hop relay channels is considered, and jamming policies that maximize distortion subject a power constraint are explored. Lastly, a regulation loss maximization subject to a soft power constraint in a fading channel is considered, where an optimal jamming policy is developed and the impact of various packet detection schemes is analyzed.
@@
Only a limited number of studies have explored the design of jamming policies using freshness metrics [22, 23, 24, 25]. Notably, Nguyen et al. [22] framed the interaction between an adversary and a scheduler as a static game by utilizing the age of information (AoI) as the performance index, and derived the optimal jamming power level for a channel represented by an $M$/$G$/$1$/$1$ queue. Xiao et al.[23] formulated such an interaction as a dynamic game by utilizing the AoI as the performance index, and derived the optimal jamming time distribution. Yang et al.[24] studied the distributed coordination of multiple sensors’ channel access behavior under jamming attacks by utilizing an AoI-based performance index, and derived optimal channel access policies when the channels are modeled as $M$/$M$/$1$, $M$/$D$/$1$, or $D$/$M$/$1$ queues. Finally, Kriouile et al. [25] considered a single-source scenario by utilizing basic AoI and AoII-based performance indices, and derived the corresponding optimal jamming policies of an omniscient adversary that knows the mismatch between the true state and the estimated state at the monitor at each time. Unlike the work in [25], we here consider an environment where the adversary does not have exact knowledge of the source states at each time. In addition, we address a multi-source scenario where the adversary selectively jams a subset of the channels.
@@ II-A Networked System Model
Our system model consists of a binary Markovian source, a remote monitor/receiver and an adversary. Time is discretized into uniform time slots. At each time slot, the Markovian source generates an update packet that is sent instantaneously over an unreliable channel to the monitor, contingent upon the state of the source. The channel’s quality, denoted as $c(t)$, is modeled as an independent and identically distributed (iid) process with two possible states: 00 and $1$. When $c(t)=1$, the packet is successfully decoded by the receiver, and when $c(t)=0$, the packet fails. The probability of successful decoding, $c(t)=1$, is given by $p$. Upon successful reception, the monitor sends an acknowledgment (ACK) back to the transmitter; otherwise, it sends a negative acknowledgment (NACK). For the Markovian source, the state transition probability, denoted by $r$, is the probability of transitioning from one state to another. We assume $r\leq 1/2$, meaning the source is more likely to remain in its current state than to switch. The adversary, acting as malicious agent within our system model, attempts to jam the channel. Let $d(t)$ denote the action of the adversary at time $t$ such that $d(t)=1$ if the adversary attempts to jam the channel, and $d(t)=0$ otherwise. When the adversary decides to jam the channel, the probability that it succeeds is equal to $q$. We consider that the adversary can intercept the ACK and NACK messages at each time.
We consider that a positive cost (a negative reward) is incurred when the adversary attempts to jam the channel due to the energy required for jamming.
@@ II-B Performance Index and It’s Dynamics
Note that, at time $t+1$, if the packet is successively transmitted, then $g(t+1)=t+1$. Therefore, at time $t+1$, the EAoII equals to $s(t+1)=s_{t+1-g(t+1)}=s_{0}$. This occurs if $c(t+1)=1$ and $d(t)=0$; or $c(t+1)=1$ and $d(t)=1$ but the adversary does not succeed in jamming the channel. However, if the packet is not successively transmitted, then $g(t+1)=g(t)$. Therefore, the EAoII will transition to $s(t+1)=s_{t+1-g(t+1)}=s_{t-g(t)+1}$. Following this observation, the transition probabilities for the EAoII are obtained as
$\Pr\big{(}s(t+1)=s_{0}|s(t)=s_{k},d(t)=0\big{)}=p,$
$\Pr\big{(}s(t+1)=s_{k+1}|s(t)=s_{k},d(t)=0\big{)}=1-p,$
$\Pr\big{(}s(t+1)=s_{0}|s(t)=s_{k},d(t)=1\big{)}=p(1-q),$
$\Pr\big{(}s(t+1)=s_{k+1}|s(t)=s_{k},d(t)=1\big{)}=1\!-p(1\!-\!q).$ (4)
@@ II-C Problem Statement
In this paper, we aim to determine the optimal jamming policy that maximizes the long-term expected average reward. Specifically, we seek to identify the sequence of actions, denoted as $\phi=(d^{\phi}(0),d^{\phi}(1),\dots)$, where $d^{\phi}(t)$ represents the jamming action taken by the attacker at time slot $t$. In this framework, $d^{\phi}(t)=1$ indicates that the adversary chooses to jam the channel at time $t$, and $d^{\phi}(t)=0$ otherwise. Taking into account the energy cost associated with jamming, the adversary’s reward at time $t$ is defined as $R(t)=s(t)-\lambda d(t)$, where $\lambda>0$ represents the jamming cost. Our objective is to solve the following optimization problem
@@ II-D Optimal Jamming Policy
In the rest of this section, we provide our theoretical results on the design of the optimal jamming policy. We begin by showing that the optimal solution is a threshold policy. Then, we provide a closed-form expression of the problem of interest, and derive the optimal threshold value as a function of $\lambda$. We reformulate our problem in (5) as infinite-horizon average-reward partially Observable Markov Decision Problem (POMDP) where $s(t)$ is the observed state at time $t$ referring to the value of EAoII, $d(t)$ the jamming action determining if the channel is under attack or not, $Pr(s\rightarrow s^{\prime}|d)$ the transition probabilities from state $s$ to $s^{\prime}$ under action $d$ that satisfy the equations given in (II-B) and $R(t)=s(t)-\lambda d(t)$ the instantaneous reward incurred at time $t$.
The optimal policy denoted $\phi^{*}$ of the problem in (5) is the one that solves the following Bellman optimality equation for each state $s=s(t)$:
$\theta+V(s)=\max_{d\in\{0,1\}}\Big{\{}s-\lambda d+\sum_{s^{\prime}\in\mathbb{S}}\Pr(s\rightarrow s^{\prime}|d)V(s^{\prime})\Big{\}}$ (6)
where $\theta$ is the optimal value of the problem, $V(s)$ is the differential reward-to-go function, and $\mathbb{S}$ is a set defined as
@@
The optimal jamming policy associated with the problem in (6) and for any given $\lambda$ is an increasing threshold policy. More specifically, there exists $n\in\mathbb{N}$ such that the prescribed action is passive, i.e., $d_{k}=0$, when $s_{k}<s_{n}$, and the prescribed action is active, i.e., $d_{k}=1$, when $s_{k}\geq s_{n}$.
@@ II-E Closed-Form Expression of the Optimal Solution
In this section, we find the explicit expression of the optimal threshold value in function of the systems parameters and $\lambda$. For that, we introduce the steady-state from of the the problem (5) as follows:
$\underset{n\in\mathbb{N}}{\text{maximize}}\ \overline{s}_{n}-\lambda\overline{d}_{n}$ (16)
where $\overline{s}_{n}$ is the average EAoII, and $\overline{d}_{n}$ is the average active attacking time (AAT) under the threshold jamming policy with parameter $s_{n}$. More specifically, we have
@@
$\overline{d}_{n}$ $=\lim_{T\to\infty}\text{inf}\>\frac{1}{T}\mathbb{E}^{n}\Big{[}\sum_{t=0}^{T-1}d(t)\big{|}s(0),\phi^{t}(n)\Big{]}$ (18)
@@
To get the expressions of $\overline{s}_{n}$ and $\overline{d}_{n}$, we will compute the stationary distribution of the discrete-time Markov chain (DTMC) that models the evolution of the EAoII under the threshold policy with parameter $s_{n}$, as illustrated in Fig. 1.
@@
$u_{n}(s_{i})=\left\{\begin{array}[]{ll}\frac{(1-q)p}{1-q+q(1-p)^{n}}&\text{if}\ i=0,\\[9.0pt] (1\!-\!p)^{i}\frac{(1-q)p}{1-q+q(1-p)^{n}}&\text{if}\ 1\leq i\leq n,\\[9.0pt] (1\!-\!p)^{n}(pq+1\!-\!p)^{i-n}&\text{if}\ i\geq n+1.\\[3.0pt] \qquad\qquad\times\frac{(1-q)p}{1-q+q(1-p)^{n}}\end{array}\right.$ (23)
@@
$\overline{s}_{n}$ $=\frac{(1-q)p}{1-q+q(1-p)^{n}}\bigg{[}\frac{r(1-p)}{p(1-y)(1-z)}$
$\quad+\frac{q(1-p)^{n}}{2pr(1-q)}-\frac{pq(1-p)^{n}(1-r)^{n+2}}{r(1-z)\big{(}1-x\big{)}}$
$\quad+\frac{pq(1-p)^{n}(1-2r)^{n+2}}{2r(1-y)\big{(}1-x\big{)}}\bigg{]}$ (25)
where $y=(1-2r)(1-p)$, $z=(1-r)(1-p)$, $w=(1-p+pq)(1-r)$ and $x=(1-p+pq)(1-2r)$
@@
Under a threshold jamming policy $s_{n}$, the average AAT $\overline{d}_{n}$ is equal to
$\overline{d}_{n}=\frac{(1-p)^{n}}{(1-q)+q(1-p)^{n}}.$ (26)
@@
Again, in the steady state, $\overline{d}_{n}=\sum_{k=n}^{\infty}u_{n}(s_{k})$. Hence, exploiting the expression of $u_{n}(\cdot)$ given in Proposition 1, we can obtain the result. ∎
Leveraging Propositions 2 and Proposition 3, we define the sequence $\lambda(s_{n})$ in the following definition, with its exact expression resembling that in (3).
@@
$\lambda(s_{n})$ is the intersection point between $\overline{s}_{n}-\lambda\overline{d}_{n}$ and $\overline{s}_{n+1}-\lambda\overline{d}_{n+1}$, i.e.,
$\lambda(s_{n})=\frac{\overline{s}_{n+1}-\overline{s}_{n}}{\overline{d}_{n+1}-\overline{d}_{n}}.$ (27)
@@
$\overline{s}_{k}\!-\!\lambda\overline{d}_{k}\geq\overline{s}_{k+1}\!-\!\lambda\overline{d}_{k+1}\ \text{for}\ \lambda\leq\lambda(s_{k}),$ (29)
$\overline{s}_{k}\!-\!\lambda\overline{d}_{k}<\overline{s}_{k+1}\!-\!\lambda\overline{d}_{k+1}\ \text{for}\ \lambda>\lambda(s_{k}),$ (30)
$\overline{s}_{n+1}\!-\!\lambda\overline{d}_{n+1}>\overline{s}_{k}\!-\!\lambda\overline{d}_{k}\ \text{for}\ \lambda>\lambda(s_{n}),\ k\leq n,$ (31)
$\overline{s}_{n+1}\!-\!\lambda\overline{d}_{n+1}\geq\overline{s}_{k}\!-\!\lambda\overline{d}_{k}\ \text{for}\ \lambda\leq\lambda(s_{n\!+\!1}),\ k>n.$ (32)